OpenStudy (kendricklamar2014):

Ohm's Law Tutorial :) (My First Tutorial) What is Ohm's Law? Ohm's Law is the relationship between the Current, the Voltage, and the Resistance. Ohm's Law states that voltage is equal to current multiplied by resistance.

OpenStudy (kendricklamar2014):

\[Ohm's ~ Law:\] \[V = IR\] \[V = Volts; ~I = Amps; ~ R = Ohms\]

OpenStudy (kendricklamar2014):

Let's try some Problems: If the voltage in a wire is 10 volts, and the current is 5 amps, what is the resistance in the wire? Solution:\[V = IR, ~10 ~Volts = 5 ~Amps * R, R = 10/5 = 2 ~Ohms\]

OpenStudy (kendricklamar2014):

What is the Voltage, when the current is 20 amps and the resistance is 7 Ohms? Solution:\[V =IR, ~V = 20 ~ Amps * 7 ~ Ohms = 140 ~ Volts\]

TheSmartOne (thesmartone):

Good job! :)

OpenStudy (kendricklamar2014):

Thanks, its still not done tho :P

OpenStudy (kendricklamar2014):

Voltage is the difference in electrical potential between two areas in an electric field. This difference in electrical potential causes current to flow. Voltage is equal to the current multiplied by the resistance in a wire.

OpenStudy (kendricklamar2014):

Quantity: Symbol: Equation(s) Standard Metric Unit: Other Units: Voltage ΔV ΔV = I*R Volt(V) J/C Current I I = V/R Amperes(A) Amp/S Power P P = ΔPE/T Watt(W) J/S Resistance R R = P*L/A Ohm(Ω) V/A Energy E or ΔPE ΔPE = ΔV*Q Joule(J) V*C

OpenStudy (kendricklamar2014):

\(\color{blue}{Created ~ By ~ KendrickLamar2014}\)

OpenStudy (alexandervonhumboldt2):

great tutorial why you hadn't tagged me to check it out? \(\Huge\color{lime}{\text{LOL}}\)

OpenStudy (kendricklamar2014):

Well I didn't tag anybody :P I totally forget. I'll do it next Tutorial :)